A car travels 25.0 km of a 50.0 km trip at an average speed of 40.0kmh^-1. It travels the second half of its journey at an average speed of 80.0kmh^-1. A truck makes the same trip but spends half of its time at an average speed of 40.0kmh^-1 and the other half of its time at an average speed of 80.0kmh^-1. Which vehicle got there in the shortest period of time?
1
Expert's answer
2020-10-25T18:26:40-0400

"time = \frac{distance}{speed}"

for the car,

total time taken "=\frac{25}{40}+\frac{25}{80}=\frac{15}{16} hours=0.9375hrs"

for the truck,

let the total time taken = "x" hours

speed "\times" time =distance

"(\frac{1}{2}x)40+(\frac{1}{2}x)80=50"

"20x+40x=50"

"60x=50"

"x=\frac{50}{60}=0.8333hrs"

the truck took the shortest time to go there.